3x+3y+6z=9
2x+y+3z=7
x+2y-z=-10

I was provided work for this problem using the matrix method. Unfortunately, I am not familiar with that method. Thanks for the help though.This is a 3-by-3 system of inequalities. In order to find the tripled pair, I need to use either the substitution of linear combination(elimination)method.
I figured out the ordered pairs, but work doesn't match up exactly.
Please help!
thank you!

Answer by Alan3354(69443)

3x+3y+6z=9 --> x+y+2z=3
2x+y+3z=7
x+2y-z=-10
----------
For this problem, use eqn 1 to eliminate x from the other 2:
x+y+2z=3 Eqn 1
x+2y-z=-10 Eqn 3
---------- Subtract
-y + 3z = 13 Eqn A
---------------------
2x+y+3z=7 Eqn 2
2x+2y+4z=6 Eqn 1 times 2
----------- Subtract
-y -z = 1 Eqn B
-y+3z = 13 Eqn A
--------------- Subtract
-4z = -12
z = 3
------
Sub for z, find y
Sub for y and z, find x